# Are We Done with MMLU?

### v2

@@ 4 Can We Fix the MMLU Dataset Automatically?
After presenting evidence of the numerous errors in the MMLU dataset, we explore the following approaches to detect these errors automatically:  [^3] The code for these experiments is available at https://github.com/aryopg/mmlu-redux.
@@
We provide the model with a straightforward instruction to classify questions into “ok” or “not ok” without introducing any demonstrations. The prompt can be found in Appendix B.
@@
Chain of Thought (CoT) prompting [15]
We encourage the model to generate reasoning steps before producing the final answer, in both zero-shot and few-shot settings. The prompt format can be found in Appendix B.
@@
Finally, we fine-tune Llama3 (8B-Instruct) model [8] using curated data and evaluate its performance on MMLU-Redux. Detailed information is provided in Appendix E.
@@ 4.1 Error Detection Experiments
To evaluate the performance of large language models (LLMs) in detecting errors in the MMLU-Redux dataset, we conduct experiments with 4 state-of-the-art models: OpenAI’s GPT-4 Turbo, GPT-4o, Anthropic’s Claude-3 Opus, and Meta’s LlamA-3-70B [8]. For each model, we test both standard prompting and Chain of Thought (CoT) prompting methods [15]. Details about the prompts are provided in Appendix B.
@@
Model Method Recall F1 Score F2 Score
@@
We consider “not ok” as the positive class, and “ok” as the negative class for the calculation of Recall, F1, and F2 scores. Based on Table 2, the Few-shot CoT setting consistently outperforms other settings across all models, suggesting that providing a small set of labelled examples along with step-by-step reasoning instructions can improve error detection performance. However, even the best-performing model, Claude-3-opus, only achieves an F2 Score of 40.29, highlighting the difficulty of this task.
Furthermore, we use retrieval-augmented prompting (RAG) to investigate the impact of external knowledge on error detection. We use BM25 to retrieve relevant paragraphs from enwiki-paragraphs and msmarco-v1-passage corpus provided by Pyserini  [17]. We retrieve the top 5 relevant paragraphs from the knowledge base using the question as the query. We then use these paragraphs as additional context in the prompt to classify the question and answer choices.
@@
Model Index Method Recall F1 Score F2 Score
@@
Based on the results presented in Table 2 and 3, we can conclude that automatic error detection in the MMLU dataset remains a challenging task, despite the availability of annotated error types. While Claude 3 Opus demonstrates the highest performance in terms of Recall, F1, and F2 Scores compared to other models, Claude 3 Opus shows the highest performance with RAG, indicating its potential for identifying errors more effectively. However, the best-performing model and method combination using RAG still achieves relatively low scores, suggesting the overall reliability of the models in detecting errors across the diverse range of subjects in MMLU is still limited. Detailed performance across all subjects can be found in Appendix C.
@@ 5 Related Work
While benchmarks often enable methodological progress, they can also be counterproductive when labelling mistakes and annotation artefacts exist. For example, Beyer et al. [18] show how annotation artefacts within the popular ImageNet benchmark [19] led to likely overstated performance gains that did not necessarily transfer to other datasets and tasks. In NLP, similar issues have been found in summarisation [20, 21] and natural language inference [22, 23, 24, 25] benchmarks.
Benchmark issues can arise from biases in the framing of the task [26]; noisy annotations [27]; web-crawls [28, 29, 30]; automated labelling processes such as crowdsourcing [31] — where annotators may be optimizing for their own incentives; human errors [32]e.g., lack of expertise on a given topic; or programmatic weak supervision [33, 34, 7], and personal [35] or group-level [36] annotator biases.
@@
The broad adoption of the MMLU benchmark for LLM evaluations [2, 1, 5, 6, 8] means identifying issues or improvements is crucial for ensuring its continued applicability. Recent studies have identified issues with labelling errors and ambiguous questions in similar benchmarks, such as MedQA [37]. Concurrent work developing the MMLU-Pro [38] benchmark also identifies a number of issues within a filtered and augmented subset of the original MMLU dataset and re-annotates parts of the dataset for inclusion in the new MMLU-Pro evaluation. However, there is currently limited literature categorising and quantifying these types of issues across the original MMLU dataset to help inform our understanding of previous results. The AI Explained YouTube Channel recently highlighted several erroneous instances in MMLU across several subjects while evaluating their SmartGPT framework in a recent popular video [12]. However, our study aims to provide a more systematic assessment of MMLU.
@@ 6 Conclusion
We analyse the Massive Multitask Language Understanding (MMLU) benchmark, driven by the necessity for rigorous evaluation of its reliability. Our analysis of 30 different MMLU subjects using a hierarchical taxonomy reveals that a significant portion of MMLU instances contain inaccuracies that could lead to misleading evaluation results. For example, 57% of the instances in the Virology and 26% in the Logical fallacies subsets were found to be inaccurate. To this end, we introduce MMLU-Redux, a thoroughly reviewed subset of the MMLU dataset [11] comprising 3,000 questions spanning the 30 MMLU subjects we analysed. The re-evaluation of LLMs using MMLU-Redux shows a significant variation in performance metrics and shifts in model rankings for several subsets, emphasising the impact that dataset quality can have on the evaluation of LLMs. Furthermore, we analyse whether it is possible to identify errors automatically; although Claude 3 Opus seems to produce the most accurate results on this task (41.9% F2 score when using retrieval-augmented generation), it is still insufficient to produce a high-quality dataset. By making MMLU-Redux publicly available, we invite the community to contribute to the endeavour of building a reliable dataset for properly evaluating the next generation of LLMs.
Limitations and Implications
Although our analysis uses a significant subset of 3,000 questions, the accuracy of this analysis, both quantitatively and qualitatively, can be further improved with additional annotation of the remaining 12,908 questions in MMLU. To properly evaluate the wide variety of subjects covered in MMLU would likely require a larger variety of experts. Therefore, we open up MMLU-Redux for additional annotation of both additional MMLU subjects and add to the already reviewed subsets. However, we acknowledge that the taxonomy we introduced to classify errors might still be prone to annotators’ personal biases. MMLU-Redux is available at https://huggingface.co/datasets/edinburgh-dawg/mmlu-redux.
Checklist
@@
      Did you describe the limitations of your work? [Yes] See Section Limitations and Implications
@@
      Did you discuss any potential negative societal impacts of your work? [No] See section Limitations and Implications
@@
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We provide details for all proposed methods. For prompting strategies, we show prompts in Appendix D For fine-tuning, we provide details about the training dataset in Appendix E.
@@
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We provide details about our resources in Appendix E
@@
      Did you include any new assets either in the supplemental material or as a URL? [Yes] MMLU-Redux is provided via a URL in the Abstract.
@@
Appendix A MMLU-Redux Error Type Statistics
@@
Table 4: Error types by subset. BQC is (1a) Bad Question Clarity, BOC is (1b) Bad Options Clarity, NCA is (2a) No Correct Answer, MCA is (2b) Multiple Correct Answers, and WGT is (2c) Wrong Ground Truth.
Subset Dataset Size OK BQC BOC NCA MCA WG
Virology 166 43 14 2 4 4 33
Logical fallacies 163 74 14 2 4 3 3
College chemistry 100 75 2 0 2 0 21
Professional law 1,534 82 4 1 0 11 2
Business ethics 100 86 14 0 0 0 0
Formal logic 126 87 1 0 2 9 1
Human aging 223 88 12 0 0 0 0
College medicine 173 96 2 0 1 0 1
High school macroeconomics 390 88 2 9 0 1 0
Global facts 100 88 1 1 4 0 6
Philosophy 311 100 0 0 0 0 0
Machine learning 112 89 3 5 0 1 2
Miscellaneous 783 90 4 1 4 1 0
Astronomy 152 94 2 2 1 0 1
Public relations 110 91 3 0 0 3 3
Professional accounting 282 94 0 1 5 0 0
Conceptual physics 235 95 0 0 0 1 4
College computer science 100 97 0 0 2 1 0
Econometrics 114 97 3 0 0 0 0
High school physics 151 97 0 0 1 0 2
High school statistics 216 98 0 0 0 0 2
Electrical engineering 145 98 1 0 0 0 1
Clinical knowledge 265 99 1 0 0 0 0
Anatomy 135 99 0 0 0 0 1
High school chemistry 203 99 1 0 0 0 0
High school mathematics 270 99 0 0 0 0 1
College mathematics 100 99 1 0 0 0 0
College physics 102 100 0 0 0 0 0
High school US history 204 100 0 0 0 0 0
High school geography 198 100 0 0 0 0 0
Total 7,263 91 3 1 1 1 3
Appendix B Prompting Methods

### v3

@@ 4 Can We Fix the MMLU Dataset Automatically?
After presenting evidence of the numerous errors in the MMLU dataset, we explore the following approaches to detect these errors automatically:
@@
We provide the model with a simple instruction to classify questions into “ok” or “not ok” without introducing any demonstrations. The prompt can be found in Appendix G.
@@
Chain of Thought (CoT) prompting
(Wei et al., 2022) We encourage the model to generate reasoning steps before producing the final answer, in both zero-shot and few-shot settings. The prompt format can be found in Appendix G.
@@
We fine-tune Llama3 8B-Instruct model (Dubey et al., 2024) using curated data and evaluate its performance on MMLU-Redux. Refer to Appendix J for more details.
@@ 4.1 Error Detection Experiments
To evaluate the performance of LLMs in detecting errors in the MMLU-Redux dataset, we conduct experiments with 4 state-of-the-art models: OpenAI’s GPT-4 Turbo, GPT-4o, Anthropic’s Claude-3 Opus, and Meta’s LlamA-3-70B (Dubey et al., 2024). For each model, we test both standard prompting and Chain of Thought (CoT) prompting methods (Wei et al., 2022). Details about the prompts are provided in Appendix G.
@@
Model Method Recall F1 F2
@@
We consider “not ok” as the positive class, and “ok” as the negative class for the calculation of Recall, F1, and F2 scores. Based on Table 3, the Few-shot CoT setting outperforms other settings across all models, suggesting that providing a small set of labelled examples with step-by-step reasoning instructions can improve error detection performance. However, even the best-performing model, Claude-3-opus, only achieves an F2 Score of 40.29, highlighting the difficulty of this task.
We then use RAG prompting to investigate the impact of external knowledge on error detection. We use BM25 to retrieve five most relevant paragraphs from enwiki-paragraphs and msmarco-v1-passage corpus provided by Pyserini  (Lin et al., 2021a) using the question as the query. We then use these paragraphs as additional context in the prompt to classify the question and answer choices.
@@
Model Index Method Recall F1 F2
@@
Based on the results presented in Table 3 and 4, we can conclude that automatic error detection in the MMLU dataset remains a challenging task, despite the availability of annotated error types. While Claude 3 Opus demonstrates the highest Recall, F1, and F2 Scores compared to other models, Claude 3 Opus shows the highest performance with RAG, indicating its potential for identifying errors more effectively. However, the best-performing model and method combination using RAG still achieves relatively low scores, suggesting the overall reliability of the models in detecting errors across the diverse range of subjects in MMLU is still limited. Detailed performance across all subjects can be found in Appendix H.
@@ 5 Related Work
While benchmarks often enable methodological progress, they can also be counterproductive when labelling mistakes and annotation artefacts exist. For example, Beyer et al. (2020) show how annotation artefacts within the popular ImageNet benchmark (Russakovsky et al., 2015) led to likely overstated performance gains that did not necessarily transfer to other datasets and tasks. In NLP, similar issues have been found in summarisation (Tejaswin et al., 2021; Bhandari et al., 2020) and natural language inference (Gururangan et al., 2018; Poliak et al., 2018; Stacey et al., 2020; Wu et al., 2022) benchmarks.
Benchmark issues can arise from biases in the framing of the task (Schwartz et al., 2017); noisy annotations (Chen et al., 2016); web-crawls (Raffel et al., 2023; Lee et al., 2021; Kaddour, 2023); automated labelling processes such as crowdsourcing (Yuen et al., 2011) — where annotators may be optimising their own incentives; human errors (Peterson et al., 2019)e.g., lack of expertise on a given topic; or programmatic weak supervision (Zhang et al., 2022; Goswami et al., 2021; Kaddour et al., 2023), and personal (Geva et al., 2019) or group-level (Liu et al., 2022) annotator biases.
@@
The broad adoption of the MMLU benchmark for LLM evaluations (Anil et al., 2023a; OpenAI, 2023; Touvron et al., 2023; Anthropic, 2024; Dubey et al., 2024) means identifying issues or improvements is crucial for ensuring its continued applicability. Recent studies have identified issues with labelling errors and ambiguous questions in similar benchmarks, such as MedQA (Saab et al., 2024). Concurrent work developing the MMLU-Pro (Wang et al., 2024) benchmark also identifies a number of issues within a filtered and augmented subset of the original MMLU dataset and re-annotates parts of the dataset for inclusion in the new MMLU-Pro evaluation. However, despite these efforts, errors from the original MMLU persist in the extended MMLU-Pro benchmark, highlighting the importance of our work in systematically identifying and addressing dataset errors. The recent Global MMLU dataset Singh et al. (2024) reevaluated MMLU to study which questions are culturally biased. We observed a similar Western bias in our qualitative analysis of the errors. Additionally, there is currently limited literature categorising and quantifying these issues across the original MMLU dataset to help inform our understanding of previous results. The AI Explained YouTube Channel recently highlighted several erroneous instances in MMLU across several subjects while evaluating their SmartGPT framework in a recent popular video (L and Stapleton, 2023). However, our study aims to provide a more systematic assessment of MMLU.
@@ 6 Conclusion
We analyse the MMLU benchmark, driven by the necessity for rigorous evaluation of its reliability. Our analysis using a hierarchical annotation protocol reveals that a significant portion of MMLU instances contain inaccuracies that could lead to misleading evaluation results. For example, 57% of the instances in the Virology and 26% in the Logical fallacies subsets were found to be inaccurate. To this end, we introduce MMLU-Redux, a thoroughly reviewed subset of the MMLU dataset (Hendrycks et al., 2021) comprising 5,700 questions spanning all 57 MMLU subjects, and estimate that MMLU has an error rate of 6.49%. The re-evaluation of LLMs using MMLU-Redux shows a significant variation in performance metrics and shifts in model rankings for several subsets, emphasising the impact that dataset quality can have on the evaluation of LLMs. Furthermore, we analyse whether it is possible to identify errors automatically; although Claude 3 Opus seems to produce the most accurate results on this task (41.9% F2 score when using retrieval-augmented generation), it is still insufficient to produce a high-quality dataset.
Limitations
Although our analysis uses a significant subset of 5,700 questions, the accuracy of this analysis, both quantitatively and qualitatively, can be further improved with additional annotation of the remaining 8,342 questions in MMLU. Therefore, we open up MMLU-Redux for additional annotation of both additional MMLU subjects and add to the already reviewed subsets. However, we acknowledge that the annotation protocol we introduced to classify errors might still be prone to the personal biases of the annotators.
Appendix A Responsible Research Checklist
@@
      Did you describe the limitations of your work? [Yes] See Section Limitations
@@
      Did you discuss any potential negative societal impacts of your work? [No] See section Limitations
@@
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We provide details for all proposed methods. For prompting strategies, we show prompts in Appendix I For fine-tuning, we provide details about the training dataset in Appendix J.
@@
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We provide details about our resources in Appendix J
@@
      Did you include any new assets either in the supplemental material or as a URL? [Yes] MMLU-Redux is provided as a supplementary material.
@@
Appendix B Data Collection and Organisation
The MMLU-Redux dataset is available at https://huggingface.co/datasets/edinburgh-dawg/mmlu-redux-2.0 (doi:10.57967/hf/2507) with a CC-BY-4.0 licence. This link also contains a public datasheet. The data used to create MMLU-Redux was obtained from cais/mmlu  [^2] https://huggingface.co/datasets/cais/mmlu, which is also utilised in the ‘lm-eval-harness’ framework (Gao et al., 2023). To ensure uniformity of our results, the language model (LM) predictions used in our performance analyses were obtained from the Holistic Evaluation of Language Models (HELM) leaderboard v1.3.0, released on May 15th, 2024.
We randomly subsampled 100 questions per MMLU subject to be presented to the annotators. The annotators are instructed to follow the introduced taxonomy by first assessing the question presentation, and then by verifying the ground truth MMLU label. The annotators were encouraged to perform an exact match search using a search engine to find occurrences of the question and multiple-choice options from credible sources. If the annotators found an exact match of the question-options pair, the annotators were asked to evaluate the answer provided by the source. Regardless of whether a label was found in the source, and whether the MMLU label is the same or not, the annotators were asked to decide whether they would follow the label using their expertise. In the cases where an exact match was not found, the annotators were asked to search for supporting evidence from trusted sources, such as government websites, textbooks, and/or other reputable organisations (e.g., World Health Organisation (WHO)). In cases where the annotators are still unsure, they were asked to annotate the question with “Expert”, denoting that the question requires more expertise. This annotated subset of MMLU is denoted as MMLU-Redux.
MMLU-Redux comprises subsampled test splits of the aforementioned thirty MMLU subsets (denoted as “config” in HF vocabulary).
Each data point in MMLU-Redux contains seven columns:
  • question (string): The original MMLU question.
  • choices (list of four strings): The original list of four choices associated with the question from the MMLU dataset.
  • answer (integer): The MMLU ground truth label in the form of an array index between 0 and 3.
  • error_type (string): The annotated error_type. The values can be one of the six error types proposed in the taxonomy (“ok”, “bad_question_clarity”, “bad_options_clarity”, “no_correct_answer”, “multiple_correct_answers”, “wrong_groundtruth”) and “expert”.
  • source (string): The potential source of the question.
  • correct_answer (string): In the case of “no_correct_answer” and “wrong_groundtruth”, the annotators can suggest the alternative correct answer.
  • potential_reason (string): A free text column for the annotators to note what they believe to have caused the error.
The question, choices, and answer columns are taken from the MMLU Hugging Face dataset (cais/mmlu).
To ensure reproducibility and facilitate further research, we will make MMLU-Redux publicly available on the Hugging Face (HF) platform in the Croissant format Akhtar et al. (2024).This format provides a standardised and easily accessible structure for the dataset, allowing researchers and practitioners to readily utilise it for their own investigations and analyses. The code to generate the data and analyses will be made available in the camera ready.
Appendix C Usage Guidelines
To reproduce our results or perform analyses similar to those presented in this study, the user may download the data and utilise all the columns. MMLU-Redux contains both correct and erroneous instances, so the user should look at the value in column “error_type” to filter samples based on the specific error type.
In those cases where the error is “no_correct_answer”, “multiple_correct_answers” or “wrong_groundtruth”, the users may examine the suggested answer reported in the “correct_answer” column. The user should consider that the questions and the options reported are the same as those in the MMLU dataset, and they have not been modified, even when affected by bad clarity.
Appendix D Maintenance Plan
D.1 General Maintenance
The datasets will be updated as needed to maintain accuracy, with announcements made for each update. These updates will be posted on the dataset’s Hugging Face page. Older versions will be preserved and documented using the git commit hash of the dataset repository.
D.2 Contributing Guidelines
We welcome contributions from the research community to enhance and expand the MMLU-Redux dataset. Primarily, we foresee two types of contribution:
Additional data annotation
Currently, MMLU-Redux covers 5,700 MMLU questions, which is a subset of the full MMLU dataset. We encourage the community to help annotate more questions to cover the entire dataset. Generally, we recommend you to follow these steps:
  1. 1.
    Familiarise yourself with the taxonomy. The taxonomy is designed to be simple and broad to cover all possible erroneous cases found in MMLU.
  2. 2.
    Open a new discussion on the Hugging Face page of MMLU-Redux (https://huggingface.co/datasets/edinburgh-dawg/MMLU-Redux/discussions/new)
  3. 3.
    Insert the title with the prefix “[ADD]” to signify additional data annotation, followed by the name of the MMLU subject you are contributing to. For example, “[ADD] Virology”.
  4. 4.
    In the description body of the issue, specify the question, the choices, the MMLU answer, the error type, the correct answer (if applicable), the question source, what you believe to be the reason for the error, and optionally an additional comment. Follow this markdown template:
    **Question**: {question}
    **Choices**: {choices} # Copy paste the python array
    **MMLU answer**: {mmlu_answer}
    **Error type**: {error_type}
    # If this field is not applicable (e.g., for bad question clarity
    # or bad options clarity), write "N/A"
    **Correct answer**: {correct_answer}
    # Ideally, a URL to where you can find the source online.
    **Source**: {source}
    # This is strictly to describe what might have caused the error
    # (e.g., incorrect HTML parsing)
    **Potential error reason**: {free text}
    # This is where the annotators can leave additional comments
    **Additional comment**: {free text}
  5. 5.
    If you intend to propose annotation of more than one question, we ask you to use a delimiter such as three-dashes (“—”).
  6. 6.
    Submit the discussion.
We aim to respond to discussions within 48 hours. The proposed changes will either be integrated into the associated MMLU subject or receive feedback for further adjustments and clarification. Discussions with fully integrated changes will be closed.
Fix to the existing annotation
The current and future annotated questions in MMLU-Redux may contain inaccuracies, particularly for questions whose answers may change over time, and other factors. We believe that the community can help keep the data up to date. We recommend everyone to propose fixes by following these steps:
  1. 1.
    Familiarise yourself with the taxonomy. The taxonomy is simple and broad to cover all possible erroneous cases found in MMLU.
  2. 2.
    Open a new discussion on the Hugging Face page of MMLU-Redux (https://huggingface.co/datasets/edinburgh-dawg/MMLU-Redux/discussions/new)
  3. 3.
    Insert the title with the prefix “[FIX]” to signify additional data annotation, followed by the name of the MMLU subject you are contributing to. For example, “[FIX] Virology”.
  4. 4.
    In the description body, specify the question, the choices, the MMLU answer, the current error type, the new error type, the current possible correct answer, the new correct answer, the question source, the potential reason that causes the error, and the optional additional comments. Follow this markdown template:
    **Question**: {question}
    **Choices**: {choices} # Copy paste the python array
    **MMLU answer**: {mmlu_answer}
    **Current Error type**: {current_error_type}
    **Current Correct answer**: {current_correct_answer}
    **New Error type**: {new_error_type}
    # If it is not applicable
    # (e.g., for bad question clarity or bad options clarity), write "N/A"
    **New Correct answer**: {new_correct_answer}
    # Ideally, a URL to where you can find the source online.
    **Source**: {source}
    # This is strictly to describe what might have caused the error
    # (e.g., incorrect HTML parsing)
    **Potential error reason**: {free text}
    # This is where the annotators can leave additional comments
    # about the data
    **Additional comment**: {free text}
  5. 5.
    If you intend to propose more than one fix, please use a delimiter such as three-dashes (“—”).
  6. 6.
    Submit the discussion.
For contributions that may not be suitable for submission via Hugging Face discussions or pull requests (e.g., bulk annotations stored in a CSV/PDF/DOCX file), we encourage you to reach out to us directly at aryo.gema@ed.ac.uk and p.minervini@ed.ac.uk, together with your HF username (if you have one and you are willing to be acknowledged). We will create a pull request which contains the proposed changes and tag your HF username. You can verify the changes and notify us via the discussion page if the pull request is ready to be merged.
D.3 Contingency Plan
We are committed to ensuring the long-term availability and accessibility of MMLU-Redux. If the primary hosting platform (i.e., Hugging Face) becomes unavailable or ceases operations, we will immediately activate our contingency measures. This includes notifying all stakeholders and users of the dataset about the temporary unavailability and redirecting them to the mirror copies available on alternative platforms (i.e., Zenodo, https://zenodo.org/records/14051942). If the primary hosting ceases operations, we will ensure the dataset is reuploaded to a new reliable hosting platform, accompanied by updated documentation and links to the new location. We will also utilise social media and other networks to inform the community about the new location of the dataset.
Appendix E Authors Statement
The authors bear all responsibility in case of violation of rights.
Appendix F MMLU-Redux Error Type Statistics
@@
Table 5: Error types by subset. BQC is (1a) Bad Question Clarity, BOC is (1b) Bad Options Clarity, NCA is (2a) No Correct Answer, MCA is (2b) Multiple Correct Answers, and WGT is (2c) Wrong Ground Truth. The percentages in the Total row are estimated via stratified sampling over the subsets.
Subset Original Size Annotated OK BQC BOC NCA MCA WGT
Virology 166 100 43 14 2 4 4 33
Logical fallacies 163 100 74 14 2 4 3 3
College chemistry 100 100 75 2 0 2 0 21
Professional law 1,534 100 82 4 1 0 11 2
Human sexuality 131 100 83 11 1 2 1 2
Business ethics 100 100 86 14 0 0 0 0
Formal logic 126 100 87 1 0 2 9 1
Human aging 223 100 88 12 0 0 0 0
Global facts 100 100 88 2 1 4 0 5
Machine learning 112 100 89 3 5 0 1 2
Miscellaneous 783 100 90 4 1 4 1 0
Abstract algebra 100 100 91 2 0 1 0 6
Public relations 110 100 91 3 0 0 3 3
High school European history 165 100 91 4 4 0 0 1
Professional accounting 282 100 94 0 1 5 0 0
Astronomy 152 100 94 2 2 1 0 1
Security studies 245 100 94 6 0 0 0 0
Conceptual physics 235 100 95 0 0 0 1 4
International law 121 100 95 0 0 0 3 2
High school biology 310 100 95 1 1 1 0 2
College medicine 173 100 96 2 0 1 0 1
Moral disputes 346 100 96 3 0 0 0 1
Marketing 234 100 96 2 2 0 0 0
Professional psychology 612 100 96 1 0 1 0 2
High school psychology 545 100 96 4 0 0 0 0
High school physics 151 100 97 0 0 1 0 2
Elementary mathematics 378 100 97 1 0 0 0 2
High school macroeconomics 390 100 97 2 0 0 1 0
College computer science 100 100 97 0 0 2 1 0
Computer security 100 100 97 0 0 0 0 3
Moral scenarios 895 100 97 3 0 0 0 0
Econometrics 114 100 97 3 0 0 0 0
Nutrition 306 100 98 1 1 0 0 0
High school microeconomics 238 100 98 1 0 0 0 1
Sociology 201 100 98 2 0 0 0 0
High school statistics 216 100 98 0 0 0 0 2
College biology 144 100 98 1 0 0 0 1
Electrical engineering 145 100 98 1 0 0 0 1
High school world history 237 100 99 0 0 0 0 1
Anatomy 135 100 99 0 0 0 0 1
High school mathematics 270 100 99 0 0 0 0 1
High school chemistry 203 100 99 1 0 0 0 0
Jurisprudence 108 100 99 0 1 0 0 0
World religions 171 100 99 1 0 0 0 0
US foreign policy 100 100 99 1 0 0 0 0
Clinical knowledge 265 100 99 1 0 0 0 0
Professional medicine 272 100 99 1 0 0 0 0
College mathematics 100 100 99 1 0 0 0 0
Management 103 100 100 0 0 0 0 0
Philosophy 311 100 100 0 0 0 0 0
Medical genetics 100 100 100 0 0 0 0 0
High school US history 204 100 100 0 0 0 0 0
High school government and politics 193 100 100 0 0 0 0 0
High school geography 198 100 100 0 0 0 0 0
High school computer science 100 100 100 0 0 0 0 0
College physics 102 100 100 0 0 0 0 0
Prehistory 324 100 100 0 0 0 0 0
Total 14,042 5,700 93.51% 2.47% 0.44% 0.62% 1.54% 1.42%
Appendix G Prompting Methods